OpenStudy (anonymous):

A cylinder has a volume of 175pi cubic units and a height of 7 units. The diameter of the cylinder is 5 units 10 units 12.5 units HELP!!!!!!!!! Because I don't understand this at ALL!!!!!!!

OpenStudy (jack1):

Volume of a cylinder = pi x r^2 x h

OpenStudy (jack1):

a diameter is twice the radius... so diameter is 2 x r \(\Large V = \pi \times r^2 \times h \) V = 175pi h = 7 pi =3.1415.... now solve for r then double it

OpenStudy (anonymous):

So would you divide everything over the the 175?

OpenStudy (jack1):

\(\Large V=π×r^2×h\) \(\Large 175\pi=π×r^2×7\) so first... divide both sides by pi then divide both sides by 7 then you should be left with r^2 = ... a number then sqrt both sides and you should be left with r then double it... and that = d

OpenStudy (anonymous):

I just got a number like 4,000 and something. I seriously don't understand :( I'm sorry...

OpenStudy (anonymous):

OOOH WAIT!

OpenStudy (anonymous):

Is the answer 10?

OpenStudy (jack1):

bam! u rock!

OpenStudy (jack1):

yep, d = 10 ;)

OpenStudy (anonymous):

Thank you so much! Sorry it took so long lol

OpenStudy (jack1):

all good dude, if it was easy it wouldn't be fun ;)